OpenStudy (please.help.me):

Can someone tell me where I went wrong in this problem? 2 cos (x/3) - Squareroot of 2=0 Solve the multiple-angle equation. (Enter your answers as a comma-separated list. Use n as an integer constant. Enter your response in radians.)

OpenStudy (please.help.me):

Screenshot with my answer that I submitted

zepdrix (zepdrix):

So when you found the solutions for x/3, what were they? pi/4 and what else?

OpenStudy (caozeyuan):

cos(x/3)=sqrt 2/2

OpenStudy (please.help.me):

pi/4 and 7pi/4

OpenStudy (please.help.me):

So \[\frac{ \pi }{ 4 } and \frac{ 7\pi }{ 4 }\]

zepdrix (zepdrix):

Mmmm ok that looks right!

OpenStudy (please.help.me):

\[\cos \frac{ x }{ 3 } =\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (please.help.me):

So what would be the final answer?

zepdrix (zepdrix):

I think you just mixed up your multiplication by 3 maybe? 3*7pi/4 = 21pi/4

zepdrix (zepdrix):

And why are you adding 5n(pi) in the first solution set? I'm really confused .. need to see your steps.

OpenStudy (please.help.me):

In this screenshot they have included an example of this problem and the answer, if it helps!

OpenStudy (please.help.me):

I am not sure what I was doing! I was going by examples from the internet trying to figure it out

zepdrix (zepdrix):

\[\large\rm \cos\left(\frac x3\right)=\frac{\sqrt2}{2}\]So within one rotation this gives us two solutions.\[\large\rm \frac x3=\frac{\pi}{4}\]\[\large\rm \frac x3=\frac{7\pi}{4}\]But we want the entire solution sets, so we need to allow for rotation,\[\large\rm \frac x3=\frac{\pi}{4}+2n \pi\]\[\large\rm \frac x3=\frac{7\pi}{4}+2n \pi\]From this point, solve for x by multiplying through by 3.

OpenStudy (please.help.me):

\[\frac{ 3\pi }{ 4 } + 6n \pi ?\]

zepdrix (zepdrix):

Ya that's one of the solution sets. Do the same for the other one.

OpenStudy (please.help.me):

21π/4+6nπ ?

zepdrix (zepdrix):

Yes

OpenStudy (please.help.me):

So those two are the final solutions?

zepdrix (zepdrix):

Yes

OpenStudy (please.help.me):

Thank you so much!!!